# Joint Transmit and Pinching Beamforming for Pinching Antenna Systems (PASS): Optimization-Based or Learning-Based?

### v2

@@ II-B Problem Formulation and Conversion
[…]
@@ III-A WMMSE-based Reformulation
$\!e_{k}\!\left(\mathbf{X},\mathbf{D},\mathbf{v}\right)\!\triangleq\!\mathbb{E}\left[\left|\widehat{y}_{k}\!-\!\widetilde{s}_{k}\right|^{2}\right]\!\!=\!\!\mathbb{E}\left[\left(v_{k}y_{k}-\widetilde{s}_{k}\right)^{H}\!\!\left(v_{k}y_{k}\!-\!\widetilde{s}_{k}\right)\right]\!\!=\!\!$
$\!\!\sum_{i\in\mathcal{K}}\!\left|v_{k}\mathbf{h}_{k}^{H}\!\!\left(\mathbf{X}\right)\!\mathbf{G}\!\!\left(\mathbf{X}\right)\!\mathbf{d}_{i}\right|^{2}\!\!+\!\!\sigma^{2}\!\left|v_{k}\right|^{2}\!\!+\!\!1\!\!-\!\!2\mathrm{Re}\!\left\{v_{k}\mathbf{h}_{k}^{H}\!\!\left(\mathbf{X}\right)\!\mathbf{G}\!\left(\mathbf{X}\right)\!\mathbf{d}_{k}\!\right\}\!,$ (11)
@@
[…]
@@
The above WMMSE-based reformulation converts the fractional SINR into a more tractable form. We could follow the BCD method to optimize (P1) if it is multi-convex with respect to (w.r.t.) each block variable $\mathbf{D}$ and $\mathbf{x}$. However, the complex exponential components and the reciprocals of distances in (III-A) render the non-convexity of (P1) w.r.t. $\mathbf{x}$, which is strongly coupled with $\mathbf{D}$. To solve this nonconvex coupled problem, we invoke the PDD theory with a majorization minimization procedure, thus ensuring the stable convergence to a critical point.
$L_{k,n,l}^{\mathrm{DC}}\left(x_{n,l}\right)\!=\!\frac{1}{2\rho}\left(u_{k,n,l}^{2}\!+\!1\right)\left(x_{n,l}\!-\!x_{k}^{\mathrm{U}}\right)^{2}\!+\!\Omega_{k,n,l}\sqrt{\left(x_{n,l}\!-\!x_{k}^{\mathrm{U}}\right)^{2}\!+\!\psi_{k,n,l}^{2}},\quad\text{(if $\Omega_{k,n,l}<0$),}$, (1a)
$\leqslant$ $\widehat{L}_{k,n,l}^{\mathrm{DC}}\left(x_{n,l}\right)=\frac{1}{2\rho}\left(u_{k,n,l}^{2}+1\right)\left(x_{n,l}-x_{k}^{\mathrm{U}}\right)^{2}+\nabla_{x_{n,l}}L_{k,n,l}^{\mathrm{CC}}\left(x_{n,l}^{(t-1)}\right)\left(x_{n,l}-x_{n,l}^{(t-1)}\right),\quad\text{(if $\Omega_{k,n,l}<0$)}.$ (1b)
$L_{k,n,l}^{\mathrm{EXP}}\left(\theta_{k,n,l}\right)\!\leqslant\!\widehat{L}_{k,n,l}^{\mathrm{EXP}}\left(\theta_{k,n,l}\right)\!\!=\!L_{k,n,l}^{\mathrm{EXP}}\left(\theta_{k,n,l}^{(t-1)}\right)+\nabla_{\theta}L_{k,n,l}^{\mathrm{EXP}}\left(\theta_{k,n,l}^{(t-1)}\right)\left(\theta_{k,n,l}-\theta_{k,n,l}^{(t-1)}\right)\!+\!\left|\frac{\varrho_{k,n,l}^{\theta}}{2}\left(\theta_{k,n,l}-\theta_{k,n,l}^{(t-1)}\right)\right|^{2},$ (1a)
$\hskip 40.00006pt\nabla_{\theta}L_{k,n,l}^{\mathrm{EXP}}\left(\theta\right)=\phi\mathrm{Re}\left\{\lambda_{k,n,l}^{u}+\frac{u_{k,n,l}s_{k,n,l}}{\rho}\right\}\sin\theta+\phi\mathrm{Im}\left\{\lambda_{k,n,l}^{u}+\frac{u_{k,n,l}s_{k,n,l}}{\rho}\right\}\cos\theta.$ (1b)
@@ III-B The Proposed MM-PDD Algorithm
We first transform (P1) into a more tractable augmented Lagrangian (AL) problem, and then establish Lipschitz gradient surrogate for convex relaxation based on the MM method. To this end, we newly introduce auxiliary variable $s_{k,n,l}=r\left(x_{n,l},\bm{\eta}_{k}^{\mathrm{U}}\right)=\sqrt{\left(x_{n,l}-x_{k}^{\mathrm{U}}\right)^{2}+\psi_{k,n,l}^{2}}$ to represent the PA-user distances, $\theta_{k,n,l}=\kappa\left(s_{k,n,l}+n_{\mathrm{eff}}x_{n,l}\right)$ to indicate the PASS-modified phase of user $k$’s signal, and $u_{k,n,l}=\sqrt{{\beta}}/\sqrt{L}e^{-i\theta_{k,n,l}}/r\left(x_{n,l},\bm{\eta}_{k}^{\mathrm{U}}\right)$ to signify the equivalent pinching beamforming coefficient of user $k$’s signal radiated by PA $a_{n,l}$. This leads to the following equality constraints:
@@
$b_{k,n,l}^{\theta}\!\triangleq\!\theta_{k,n,l}\!-\!\kappa\left(s_{k,n,l}+\!n_{\mathrm{eff}}x_{n,l}\right)\!=\!0,$ (1b)
$b_{k,n,l}^{s}\!\triangleq\!\sqrt{\left(x_{n,l}\!-\!x_{k}^{\mathrm{U}}\right)^{2}\!+\!\psi_{k,n,l}^{2}}\!-\!s_{k,n,l}\!=\!0,$ (1c)
where $\phi\triangleq\sqrt{\beta}/\sqrt{L}$ is a constant, and $\psi_{k,n,l}=\sqrt{\left(y_{n,l}^{\mathrm{PA}}-y_{k}^{u}\right)^{2}+\left(h^{\mathrm{PA}}\right)^{2}}$ is fixed for each channel realization. Moreover, we define $q_{k,k^{\prime}}=\mathbf{h}_{k}^{H}\left(\mathbf{X}\right)\mathbf{G}\left(\mathbf{X}\right)\mathbf{d}_{k^{\prime}}$ as the path response of user $k$ to receive the signal of user $k^{\prime}$. From the above definitions, $q_{k,k^{\prime}}$ satisfies the equality constraint:
$b_{k,k^{\prime}}^{q}\triangleq q_{k,k^{\prime}}-\mathbf{u}_{k}^{T}\bm{\Sigma}\mathbf{d}_{k^{\prime}}=0,~{}\forall{k,k^{\prime}}\in\mathcal{K},$ (2)
where $\mathbf{u}_{k}^{T}=\left[\mathbf{u}_{k,1}^{T},\mathbf{u}_{k,2}^{T},\dots,\mathbf{u}_{k,N}^{T}\right]\in\mathbb{C}^{1\times M}$ stacks the pinching beamforming coefficients for all the PAs, and $\mathbf{u}_{k,n}^{T}=\left[u_{k,n,1},u_{k,n,2},\dots,u_{k,n,L}\right]^{T}\in\mathbb{C}^{1\times L}$ is the pinching beamforming vector of user $k$ using PAs in subset $\mathcal{L}_{n}$. Moreover, $\bm{\Sigma}=\mathrm{blkdiag}\left(\mathbf{1}_{L\times 1},\mathbf{1}_{L\times 1},\dots,\mathbf{1}_{L\times 1}\right)\in\mathbb{R}^{M\times N}$ is a block diagonal matrix.
@@
[…]
$\!\!\sum_{k\in\mathcal{K}}\!\!\Bigg{(}\!\!\left\|\mathbf{B}_{k}^{u}\!\!+\!\!\rho\bm{\lambda}_{k}^{u}\right\|^{2}\!+\!\left\|\mathbf{B}_{k}^{\theta}\!\!+\!\!\rho\bm{\lambda}_{k}^{\theta}\right\|^{2}\!+\!\left\|\mathbf{B}_{k}^{s}\!\!+\!\!\rho\bm{\lambda}_{k}^{s}\right\|^{2}\!+\!\left\|\mathbf{b}_{k}^{q}\!\!+\!\!\rho\bm{\lambda}_{k}^{q}\right\|^{2}\!\!\Bigg{)},$
[…]
where $\mathbf{Q}=\left\{q_{k,k^{\prime}}\right\}\in\mathbb{C}^{K\times K}$ and $\mathbf{U}=\left[\mathbf{u}_{1},\mathbf{u}_{2},\dots,\mathbf{u}_{K}\right]\in\mathbb{R}^{M\times K}$. $\mathbf{B}_{k}^{u}=\left\{b_{k,n,l}^{u}\right\}$, $\mathbf{B}_{k}^{\theta}=\left\{b_{k,n,l}^{\theta}\right\}$, and $\mathbf{B}_{k}^{s}=\left\{b_{k,n,l}^{s}\right\}$ stack the residuals of the equality constraints in (1), and $\mathbf{b}_{k}^{q}=\left\{b_{k^{\prime},k}^{q}\right\}$ stacks the residual of equality constraint (2). Moreover, $\bm{\lambda}_{k}^{u}\in\mathbb{C}^{N\times L}$, $\bm{\lambda}_{k}^{\theta}\in\mathbb{C}^{N\times L}$, $\bm{\lambda}_{k}^{s}\in\mathbb{R}^{N\times L}$, and $\bm{\lambda}_{k}^{q}\in\mathbb{C}^{K\times 1}$ are dual variables corresponding to (1) and (2), respectively. By fixing $v_{k}^{opt}$, the MSE is rewritten as a convex function of w.r.t. $q_{k,k^{\prime}}$:
@@
Now, the nonconvexity of problem (P2) only lies in the AL term $\frac{1}{2\rho}\left\|\mathbf{B}_{k}^{u}+\rho\bm{\lambda}_{k}^{u}\right\|^{2}+\frac{1}{2\rho}\left\|\mathbf{B}_{k}^{s}+\rho\bm{\lambda}_{k}^{s}\right\|^{2}$ w.r.t. $\mathbf{x}$ and $\bm{\theta}$, which is the sum of terms
$L_{k,n,l}^{\mathrm{NC}}$ $\!=\!\!\frac{1}{2\rho}\!\left(\!u_{k,n,l}\!\sqrt{\!\left(x_{n,l}\!\!-\!\!x_{k}^{\mathrm{U}}\right)^{2}\!\!+\!\!\psi_{k,n,l}^{2}}\!\!-\!\!\phi e^{-i\theta_{k,n,l}}\!\!+\!\rho\lambda_{k,n,l}^{u}\right)^{2}$
$\!+\!\frac{1}{2\rho}\!\left(\!\sqrt{\!\left(x_{n,l}\!-\!x_{k}^{\mathrm{U}}\right)^{2}\!+\!\psi_{k,n,l}^{2}}\!\!-\!\!s_{k,n,l}\!+\!\rho\lambda_{k,n,l}^{s}\!\right)^{2}\!\!\!.$ (4)
@@
On the one hand, $L_{k,n,l}^{\mathrm{NC}}$ is a difference of convex (D.C.) function w.r.t. $x_{n,l}$. The D.C. function $L_{k,n,l}^{\mathrm{DC}}\left(x_{n,l}\right)$ is defined in (1a), as shown in the bottom of the next page, where $\Omega_{k,n,l}\triangleq\mathrm{Re}\big{\{}u_{k,n,l}^{H}\big{(}\lambda_{k,n,l}^{u}-\frac{\phi e^{-i\theta_{k,n,l}}}{\rho}\big{)}\big{\}}+\big{(}\lambda_{k,n,l}^{s}-\frac{s_{k,n,l}}{\rho}\big{)}$ is the coefficient of the convex function $\sqrt{\left(x_{n,l}-x_{k}^{\mathrm{U}}\right)^{2}+\psi_{k,n,l}^{2}}$. When $\Omega_{k,n,l}$ is a negative coefficient, function $L_{k,n,l}^{\mathrm{DC}}\left(x_{n,l}\right)$ contains a concave component $L^{\mathrm{CC}}\!\left(x_{n,l}\right)=\Omega_{k,n,l}\sqrt{\left(x_{n,l}-x_{k}^{\mathrm{U}}\right)^{2}+\psi_{k,n,l}^{2}}$. We handle the D.C. component using the concave-convex procedure (CCCP), which is a special case of MM for D.C. programming [16, 17]. CCCP utilizes the first-order Taylor expansion of $L^{\mathrm{CC}}\!\left(x_{n,l}\right)$ as its tight upper bound, as presented in (1) in the bottom of this page, where the derivative $\nabla_{x}L_{k,n,l}^{\mathrm{CC}}\!\left(x_{n,l}^{(t-1)}\right)$ can be given by
$\nabla_{x_{n,l}}L_{k,n,l}^{\mathrm{CC}}\left(x_{n,l}^{(t-1)}\right)\!=\!\frac{\Omega_{k,n,l}\left(x_{n,l}^{(t-1)}-x_{k}^{\mathrm{U}}\right)}{\left(x_{n,l}^{(t-1)}-x_{k}^{\mathrm{U}}\right)^{2}\!+\!\psi_{k,n,l}^{2}}.$ (1)
On the other hand, $L_{k,n,l}^{\mathrm{NC}}$ involves the complex exponential component $L^{\mathrm{EXP}}\left(\theta_{k,n,l}\right)=-\phi\mathrm{Re}\big{\{}\big{(}\lambda_{k,n,l}^{u}+\frac{u_{k,n,l}s_{k,n,l}}{\rho}\big{)}e^{i\theta_{k,n,l}}\big{\}}$, which periodically switch between convex and concave functions over the feasible range of $\theta_{k,n,l}\in\mathbb{R}$. To convert this intractable nonconvex component, we construct Lipschitz gradient surrogate based on the MM method [18], which can be defined as follows.
@@
$\mathcal{S}\left(\bm{\theta},\bm{\theta}^{(t-1)}\right)$ $=f\left(\bm{\theta}^{(t-1)}\right)+$
$\!\!\left(\!\nabla_{\bm{\theta}}f\!\left(\bm{\theta}^{(t-1)}\!\right)\!\right)^{\!T}\!\!\!\!\left(\bm{\theta}\!-\!\bm{\theta}^{(t-1)}\!\right)\!\!+\!\!\frac{\varrho}{2}\left\|\bm{\theta}\!-\!\bm{\theta}^{(t-1)}\!\right\|^{2}\!,$ (2)
@@
The Lipschitz gradient surrogate $\mathcal{S}\left(\bm{\theta},\bm{\theta}^{(t-1)}\right)$ establishes an MM upper bound of $f\left(\bm{\theta}\right)$ when the gradient $\nabla_{\bm{\theta}}f\left(\bm{\theta}^{(t-1)}\!\right)$ is $\varrho$-Lipschitz continuous [18, 17]. Compared to the first-order Taylor expansion, Lipschitz gradient surrogate exploits an extra second-order regularization term, which achieves a strict convexity approximation and thus improves the convergence guarantees. To apply MM, we introduce the following lemma to construct Lipschitz gradient surrogate of the exponential term $L^{\mathrm{EXP}}\left(\theta_{k,n,l}\right)=-\phi\mathrm{Re}\left\{\left(\lambda_{k,n,l}^{u}+\frac{u_{k,n,l}s_{k,n,l}}{\rho}\right)e^{i\theta_{k,n,l}}\right\}$.

### v3

@@ II-B Problem Formulation and Conversion
s.t. $\text{(C1): }x_{n,l}\!-\!x_{n,l-1}\geqslant\Delta_{\min},~\forall 1<l\leqslant L,\forall n\in\mathcal{N},$ (10b)
$\text{(C2): }0\leqslant x_{n,l}\leqslant S_{\mathrm{x}},~\forall l\in\mathcal{M},$ (10c)
$\text{(C3): }\sum_{k\in\mathcal{K}}\left\|\mathbf{d}_{k}\right\|_{2}^{2}\leqslant P,$ (10d)
@@ III-A WMMSE-based Reformulation
$\hskip-5.0pt\!e_{k}\!\left(\mathbf{X},\mathbf{D},\mathbf{v}\right)\!\triangleq\!\mathbb{E}\left[\left|\widehat{y}_{k}\!-\!\widetilde{s}_{k}\right|^{2}\right]\!\!=\!\!\mathbb{E}\left[\left(v_{k}y_{k}-\widetilde{s}_{k}\right)^{H}\!\!\left(v_{k}y_{k}\!-\!\widetilde{s}_{k}\right)\right]\!\!=\!\!\\ \!\!\sum_{i\in\mathcal{K}}\!\left|v_{k}\mathbf{h}_{k}^{H}\!\!\left(\mathbf{X}\right)\!\mathbf{G}\!\!\left(\mathbf{X}\right)\!\mathbf{d}_{i}\right|^{2}\!+\sigma^{2}\!\left|v_{k}\right|^{2}\!+\!1\!-\!2\mathrm{Re}\!\left\{v_{k}\mathbf{h}_{k}^{H}\!\!\left(\mathbf{X}\right)\!\mathbf{G}\!\left(\mathbf{X}\right)\!\mathbf{d}_{k}\!\right\}\!,$ (11)
@@
$\text{(P1)}~\min_{\mathbf{X},\mathbf{D},\mathbf{v},\bm{\alpha}}$ $\sum_{k=1}^{K}\left(\alpha_{k}e_{k}\left(\mathbf{X},\mathbf{D},v_{k}\right)-\log_{2}\alpha_{k}\right),$ (12a)
s.t. $\text{(C1) - (C3)},$ (12b)
@@
The above WMMSE-based reformulation converts the fractional SINR into a more tractable form. We could follow the BCD method to optimize (P1) if it is multi-convex with respect to (w.r.t.) each block variable $\mathbf{D}$ and $\mathbf{x}$. However, the complex exponential components and the reciprocals of distances in (11) render the non-convexity of (P1) w.r.t. $\mathbf{x}$, which is strongly coupled with $\mathbf{D}$. To solve this nonconvex coupled problem, we invoke the PDD theory with a majorization minimization procedure, thus ensuring the stable convergence to a critical point.
[…]
$\leqslant\!L_{k,n,l}^{\mathrm{MM}}\left(x_{n,l}\right)L_{k,n,l}^{\mathrm{conv}}\!\left(x_{n,l}\right)\!\triangleq\!\!+\!\widehat{L}_{k,n,l}^{\!\mathrm{CC}}\!\left(x_{n,l}\right)\!+\!\widehat{L}_{k,n,l}^{\mathrm{NC}}\!\left(x_{n,l}\right),$ (1b)
$\widehat{L}_{k,n,l}^{\!\mathrm{CC}}\!\left(x_{n,l}\right)\!$ $=\!L_{k,n,l}^{\!\mathrm{CC}}\!\left(x_{n,l}^{(t-1)}\right)\!\!+\!\nabla_{\!x}L_{k,n,l}^{\mathrm{CC}}\!\!\left(x_{n,l}^{(t-1)}\right)\!\!\left(x_{n,l}\!-\!x_{n,l}^{(t-1)}\!\right)\!\!=\!L_{k,n,l}^{\!\mathrm{CC}}\!\left(x_{n,l}^{(t-1)}\right)\!\!+\!\tfrac{\Omega_{k,n,l}\left(x_{n,l}^{(t-1)}\!-\!x_{k}^{\mathrm{U}}\right)}{\left(x_{n,l}^{(t-1)}-x_{k}^{\mathrm{U}}\right)^{2}\!+\!\psi_{k,n,l}^{2}}\!\!\left(x_{n,l}\!-\!x_{n,l}^{(t-1)}\!\right)\!,$ (1c)
$\widehat{L}_{k,n,l}^{\mathrm{NC}}\!\left(x_{n,l}\right)\!$ $\triangleq\!\tfrac{x_{n,l}^{3}\!+\!\left(x_{k}^{\mathrm{U}}\right)^{2}\!+\!\big(2\psi_{k,n,l}^{2}\!+\!\left(x_{n,l}^{(t-1)}\!-\!x_{k}^{\mathrm{U}}\right)^{2}\!-\!4x_{k}^{\mathrm{U}}x_{n,l}^{(t-1)}\big)x_{n,l}\!+\!2x_{k}^{\mathrm{U}}\left(x_{n,l}^{(t-1)}\right)^{2}}{2\sqrt{\left(x_{n,l}^{(t-1)}-x_{k}^{\mathrm{U}}\right)^{2}+\psi_{k,n,l}^{2}}}\!\geqslant\!L_{k,n,l}^{\mathrm{NC}}\!\left(x_{n,l}\right)\!\triangleq\!x_{n,l}\sqrt{\left(x_{n,l}\!-\!x_{k}^{\mathrm{U}}\right)^{2}\!+\!\psi_{k,n,l}^{2}}.$ (1d)
$L_{k,n,l}^{\mathrm{ex}}\left(\theta_{k,n,l}\right)\!\leqslant\!\widehat{L}_{k,n,l}^{\mathrm{ex}}\left(\theta_{k,n,l}\right)\!\!=\!L_{k,n,l}^{\mathrm{ex}}\left(\theta_{k,n,l}^{(t-1)}\right)+\nabla_{\theta}L_{k,n,l}^{\mathrm{ex}}\left(\theta_{k,n,l}^{(t-1)}\right)\left(\theta_{k,n,l}-\theta_{k,n,l}^{(t-1)}\right)\!+\!\left|\tfrac{\varrho_{k,n,l}^{\theta}}{2}\left(\theta_{k,n,l}-\theta_{k,n,l}^{(t-1)}\right)\right|^{2},$ (1a)
$\nabla_{\theta}L_{k,n,l}^{\mathrm{ex}}\left(\theta\right)\!=\!\phi\mathrm{Re}\left\{\lambda_{k,n,l}^{u}+\tfrac{u_{k,n,l}r_{k,n,l}}{\rho}\right\}\sin\theta+\phi\mathrm{Im}\left\{\lambda_{k,n,l}^{u}+\tfrac{u_{k,n,l}r_{k,n,l}}{\rho}\right\}\cos\theta.$ (1b)
@@ III-B The Proposed MM-PDD Algorithm
We first transform (P1) into a more tractable augmented Lagrangian (AL) problem, and then establish Lipschitz gradient surrogate for convex relaxation based on the MM method. To this end, we newly introduce auxiliary variable $\theta_{k,n,l}=\kappa\left(r\left(x_{n,l},\bm{\eta}_{k}^{\mathrm{U}}\right)+n_{\mathrm{eff}}x_{n,l}\right)$ to indicate the PASS-modified phase of user $k$’s signal, and $u_{k,n,l}=\phi e^{-i\theta_{k,n,l}}/r\left(x_{n,l},\bm{\eta}_{k}^{\mathrm{U}}\right)$ to signify the equivalent pinching beamforming coefficient of user $k$’s signal radiated by PA $a_{n,l}$, where $\phi\triangleq\sqrt{{\beta}}/\sqrt{L}$ is a constant. This leads to the following equality constraints:
@@
$b_{k,n,l}^{\theta}\!\triangleq\!\theta_{k,n,l}\!-\!\kappa\!\left(\!\sqrt{\left(x_{n,l}\!-\!x_{k}^{\mathrm{U}}\!\right)^{2}\!+\!\psi_{k,n,l}^{2}}\!+\!n_{\mathrm{eff}}x_{n,l}\!\right)\!=\!0,$ (1b)
where $\psi_{k,n,l}\!\triangleq\!\sqrt{\left(y_{n,l}^{\mathrm{PA}}\!-\!y_{k}^{u}\right)^{2}\!+\!\left(h^{\mathrm{PA}}\right)^{2}}$ is fixed for each channel realization. Moreover, we define auxiliary variable $q_{k,k^{\prime}}=\mathbf{h}_{k}^{H}\left(\mathbf{X}\right)\mathbf{G}\left(\mathbf{X}\right)\mathbf{d}_{k^{\prime}}$ as the path response of user $k$ to receive the signal of user $k^{\prime}$. From the above definitions, $q_{k,k^{\prime}}$ satisfies the equality constraint:
$b_{k,k^{\prime}}^{q}\triangleq q_{k,k^{\prime}}-\mathbf{u}_{k}^{\top}\bm{\Sigma}\mathbf{d}_{k^{\prime}}=0,~\forall{k,k^{\prime}}\in\mathcal{K},$ (2)
where $\mathbf{u}_{k}^{\top}=\left[\mathbf{u}_{k,1}^{\top},\mathbf{u}_{k,2}^{\top},\dots,\mathbf{u}_{k,N}^{\top}\right]\in\mathbb{C}^{1\times M}$ stacks the pinching beamforming coefficients for all the PAs, and $\mathbf{u}_{k,n}^{\top}=\left[u_{k,n,1},u_{k,n,2},\dots,u_{k,n,L}\right]\in\mathbb{C}^{1\times L}$ is the pinching beamforming vector of user $k$ using PAs in subset $\mathcal{L}_{n}$. Moreover, $\bm{\Sigma}=\mathrm{blkdiag}\left(\mathbf{1}_{L\times 1},\mathbf{1}_{L\times 1},\dots,\mathbf{1}_{L\times 1}\right)\in\mathbb{R}^{M\times N}$ is a block diagonal matrix.
@@
$\text{(P2)}~\min_{\mathbf{x},\mathbf{D},\bm{\alpha},\mathbf{v},\mathbf{S},\mathbf{U},\bm{\theta},\mathbf{Q}}$~ $\sum_{k=1}^{K}\!\left(\alpha_{k}e_{k}\!\left(\mathbf{Q},v_{k}^{opt}\right)\!\!-\!\log_{2}\!\alpha_{k}\right)\!$
$\hskip-60.00009pt\!\!+\!\frac{1}{2\rho}\!\sum_{k\in\mathcal{K}}\!\!\Bigg(\!\!\left\|\mathbf{B}_{k}^{u}\!\!+\!\!\rho\bm{\lambda}_{k}^{u}\right\|^{2}\!+\!\left\|\mathbf{B}_{k}^{\theta}\!\!+\!\!\rho\bm{\lambda}_{k}^{\theta}\right\|^{2}\!+\!\left\|\mathbf{b}_{k}^{q}\!\!+\!\!\rho\bm{\lambda}_{k}^{q}\right\|^{2}\!\!\Bigg),$
$\hskip-40.00006pt\text{s.t.}~\text{(C1) - (C3)},$
where $\mathbf{Q}=\left\{q_{k,k^{\prime}}\right\}\in\mathbb{C}^{K\times K}$ and $\mathbf{U}=\left[\mathbf{u}_{1},\mathbf{u}_{2},\dots,\mathbf{u}_{K}\right]\in\mathbb{R}^{M\times K}$. $\mathbf{B}_{k}^{u}=\left\{b_{k,n,l}^{u}\right\}$ and $\mathbf{B}_{k}^{\theta}=\left\{b_{k,n,l}^{\theta}\right\}$, stack the residuals of the equality constraints in (1), and $\mathbf{b}_{k}^{q}=\left\{b_{k^{\prime},k}^{q}\right\}$ stacks the residuals of equality constraint (2). Moreover, $\bm{\lambda}_{k}^{u}\in\mathbb{C}^{N\times L}$, $\bm{\lambda}_{k}^{\theta}\in\mathbb{C}^{N\times L}$, and $\bm{\lambda}_{k}^{q}\in\mathbb{C}^{K\times 1}$ are dual variables corresponding to (1) and (2), respectively. By fixing $v_{k}^{opt}$, the MSE is rewritten as a convex function of w.r.t. $q_{k,k^{\prime}}$:
@@
Now, the nonconvexity of problem (P2) only lies in the AL term $\left\|\mathbf{B}_{k}^{u}+\rho\bm{\lambda}_{k}^{u}\right\|^{2}+\left\|\mathbf{B}_{k}^{\theta}+\rho\bm{\lambda}_{k}^{\theta}\right\|^{2}$ w.r.t. $\mathbf{x}$ and $\bm{\theta}$, which is the sum of terms
$\begin{split}&L_{k,n,l}\!=\!\!\left|u_{k,n,l}\sqrt{\!\left(x_{n,l}\!-\!x_{k}^{\mathrm{U}}\right)^{2}\!\!+\!\psi_{k,n,l}^{2}}\!-\!\!\phi e^{-i\theta_{k,n,l}}\!\!+\!\rho\lambda_{k,n,l}^{u}\right|^{2}\!+\!\\ &\!\left(\!\!\theta_{k,n,l}\!-\!\kappa\!\sqrt{\!\left(x_{n,l}\!-\!x_{k}^{\mathrm{U}}\right)^{2}\!+\!\psi_{k,n,l}^{2}}\!-\!\kappa n_{\mathrm{eff}}x_{n,l}\!+\!\rho\lambda_{k,n,l}^{\theta}\!\!\right)^{\!\!2}\!.\end{split}$ (4)
@@
Regarding $x_{n,l}$, the AL function (4) can be rewritten as function $L_{k,n,l}^{\mathrm{AL}}\left(x_{n,l}\right)$ defined in (1a) in the bottom of the next page, which consists of a difference of convex (D.C.) function $L_{k,n,l}^{\mathrm{DC}}\big(x_{n,l}\big)$, and a nonconvex component $L_{k,n,l}^{\mathrm{NC}}\big(x_{n,l}\big)\!\triangleq\!\frac{\kappa^{2}n_{\mathrm{eff}}}{\rho}x_{n,l}\sqrt{\big(x_{n,l}\!-\!x_{k}^{\mathrm{U}}\big)^{2}+\psi_{k,n,l}^{2}}$. We construct the MM surrogate of $L_{k,n,l}^{\mathrm{AL}}\left(x_{n,l}\right)$ in (1a) as follows. First, for the D.C. function $L_{k,n,l}^{\mathrm{DC}}\big(x_{n,l}\big)$, we can observe that $L_{k,n,l}^{\mathrm{conv}}\big(x_{n,l}\big)$ is convex over $x_{n,l}$, and $L^{\mathrm{CC}}\!\left(x_{n,l}\right)\!\triangleq\!\Omega_{k,n,l}\sqrt{\left(x_{n,l}\!-\!x_{k}^{\mathrm{U}}\right)^{2}\!+\!\psi_{k,n,l}^{2}}$ is a concave component when the coefficient $\Omega_{k,n,l}\triangleq\mathrm{Re}\big\{u_{k,n,l}^{H}\big(\lambda_{k,n,l}^{u}\!-\!\tfrac{\phi e^{-i\theta_{k,n,l}}}{\rho}\big)\big\}\!+\!\big(\lambda_{k,n,l}^{s}\!-\!\tfrac{r_{k,n,l}}{\rho}\big)$ is negative. We handle the D.C. component using the concave-convex procedure (CCCP), which is a special case of MM for D.C. programming [16, 17]. CCCP utilizes the first-order Taylor expansion $\widehat{L}^{\mathrm{CC}}\!\left(x_{n,l}\right)$ as the tight upper bound of the concave component $L^{\mathrm{CC}}\!\left(x_{n,l}\right)$ if $\Omega_{k,n,l}\leqslant 0$, as presented in (1c) in the bottom of this page. Secondly, for the nonconvex component $L_{k,n,l}^{\mathrm{NC}}\big(x_{n,l}\big)$, from Jensen’s inequality we have $\sqrt{x}\leqslant\sqrt{x^{(t-1)}}+\frac{x-x^{(t-1)}}{2\sqrt{x^{(t-1)}}}\leqslant\frac{x+x^{(t-1)}}{2\sqrt{x^{(t-1)}}}$. Applying the above inequality we have
$\sqrt{\!\left(x_{n,l}\!-\!x_{k}^{\mathrm{U}}\right)^{2}\!+\!\psi_{k,n,l}^{2}}\!\leqslant\!\tfrac{\left(x_{n,l}\!-\!x_{k}^{\mathrm{U}}\right)^{2}\!\!+\!\left(x_{n,l}^{\!(t-1)}\!-\!x_{k}^{\mathrm{U}}\right)^{2}\!\!+\!2\psi_{k,n,l}^{2}}{2\sqrt{\!\left(x_{n,l}^{(t-1)}\!-\!x_{k}^{\mathrm{U}}\right)^{2}\!+\!\psi_{k,n,l}^{2}}}\!.$ (1)
Substituting (1) into $L_{k,n,l}^{\mathrm{NC}}\big(x_{n,l}\big)$ and exploiting the first-order Taylor expansion $x_{n,l}^{2}\!\leqslant\!-\!\big(x_{n,l}^{(t-1)}\big)^{2}\!-\!2x_{n,l}^{(t-1)}\big(x_{n,l}-x_{n,l}^{(t-1)}\big)\!=\!\big(x_{n,l}^{(t-1)}\big)^{2}-2x_{n,l}^{(t-1)}x_{n,l}$, the MM surrogate of $L_{k,n,l}^{\mathrm{NC}}\big(x_{n,l}\big)$ can be derived, as shown in (1d).
Regarding $\theta_{k,n,l}$, given a fixed PA-user distance $r_{k,n,l}=r(x_{n,l},\bm{\eta}_{k}^{\mathrm{U}})$, the objective function $L_{k,n,l}^{\mathrm{NC}}$ involves the complex exponential component $L^{\mathrm{ex}}\left(\theta_{k,n,l}\right)\!=\!-\phi\mathrm{Re}\big\{\big(\lambda_{k,n,l}^{u}\!+\!\frac{u_{k,n,l}r_{k,n,l}}{\rho}\big)e^{i\theta_{k,n,l}}\big\}$, which periodically switch between convex and concave functions over the feasible range of $\theta_{k,n,l}\in\mathbb{R}$. To convert this intractable nonconvex component, we construct Lipschitz gradient surrogate based on the MM method [18], which can be defined as follows.
@@
$\mathcal{S}\left(\bm{\theta},\bm{\theta}^{(t-1)}\right)=f\left(\bm{\theta}^{(t-1)}\right)+\\ \!\!\left(\!\nabla_{\bm{\theta}}f\!\left(\bm{\theta}^{(t-1)}\!\right)\!\right)^{\!T}\!\!\!\!\left(\bm{\theta}\!-\!\bm{\theta}^{(t-1)}\!\right)\!\!+\!\!\frac{\varrho}{2}\left\|\bm{\theta}\!-\!\bm{\theta}^{(t-1)}\!\right\|^{2}\!,$ (2)
@@
The Lipschitz gradient surrogate $\mathcal{S}\left(\bm{\theta},\bm{\theta}^{(t-1)}\right)$ establishes an MM upper bound of $f\left(\bm{\theta}\right)$ when the gradient $\nabla_{\bm{\theta}}f\left(\bm{\theta}^{(t-1)}\!\right)$ is $\varrho$-Lipschitz continuous [18, 17]. Compared to the first-order Taylor expansion, Lipschitz gradient surrogate exploits an extra second-order regularization term, which achieves a strict convexity approximation and thus improves the convergence guarantees. To apply MM, we introduce the following lemma to construct Lipschitz gradient surrogate of the exponential term $L^{\mathrm{ex}}\left(\theta_{k,n,l}\right)=-\phi\mathrm{Re}\left\{\left(\lambda_{k,n,l}^{u}+\frac{u_{k,n,l}r_{k,n,l}}{\rho}\right)e^{i\theta_{k,n,l}}\right\}$.